3 dices are rolled together what is the probability of getting different nos on the 3 dices???????
in what respect does this ques is diff from prev one?
give the solutions also>>>>>>>>>>

Grade:12

2 Answers

jitender lakhanpal
12 years ago

Dear Nivi,

sample space = 63

and the number of favourable cases can be visualised as arranging 6 different objects in 3 different places 

this is same as the definition of permutation so the number of favourable cases is P(6,3)=6*5*4

so the probability = (6*5*4)/216 = 5/9

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

jitender

 

New layer...
ayush jain
12 years ago

n(e)=6*5*4=120

n(s)=6*6*6=216

than probility=120/216=5/9.....................(ans)